Let me tell you why I am Using this site, Today I was taking my test and had no CLUEE on what to do or how, So today I looked up lessons and problems to help my self out, in order to atleasst finish my exam tomorow and get a somewhat good grade! so As I was doing my lesson I came across some problems that I couldn't solve properly. Please answer. Thanks for your helpp! ( any numbers with a / sign is a fraction) Please Show all work so I under stand how the problem is done, Just a answer wouldn't help if I didn't know how the problem was solve. Also sorry if this is in the wrong section.+
#1: 5-9w=23 #2: 3/2a-8=7 #3: 7=c/-5+3 #4: 6+-12+h/-7

Answer by stanbon(57323)

#1: 5-9w=23
Add 9w to both sides
Subtract 23 from both sides
---
9w = -18
Divide both sides by 9 to get:
w = -2
===================================
#2: (3/2)a-8 = 7
Add 8 to both sides:
(3/2)a = 15
---
Multiply both sides by (2/3) to get:
(2/3)(3/2)a (2/3)15
a = 10
========================================
#3: 7=c/-5+3
7 = (c/-5) + 3
Subtract 3 from both sides
4 = (c/-5)
Multiply both sides by -5 to get:
c = -5*4
c = -20
========================================
#4: 6+-12+h/-7
Not an equation.
=========================
Comment: It would help if you used parentheses when you post your problems.
=====================
Cheers,
stan H.
==============